What is the charge of a body whose p A body with an electron of 15 x 10 -20 has a charge of 6 x 10 -10 . Determine the number of protons the body has?roton is 4 x 10 15 and 5 x 10 15 ?
1
Expert's answer
2020-12-14T07:16:21-0500

The charge from electrons of a body with mass of electrons M=15x10-20 kg can be found if we find the number of electrons. The number of electrons is the mass of electrons divided by mass of a single electron me:


"N_e=\frac{M}{m_e}."

The charge of a body is


"q=N_ee=\frac{M}{m_e}e,\\\space\\
q=\frac{15\cdot10^{-20}}{9.1\cdot10^{-31}}\cdot1.609\cdot10^{-19}=26.4\cdot10^{-9}\text{ C},"

or 26.4 nC.


Electrically neutral body has equal number of electrons and protons;


"N_p=N_e=\frac{M}{m_e}=1.65\cdot10^{11}."